OpenStudy (anonymous):

Find the slope of the line tangent to the curve at the given value of x f(x)=x^2+5x at x=-6

OpenStudy (anonymous):

f'(x)=2x+5 The question is to find f'(-6), so plug in -6 above.

OpenStudy (anonymous):

-66? That seems off

OpenStudy (anonymous):

2*(-6)+5

OpenStudy (anonymous):

isnt it plus 5x?

OpenStudy (anonymous):

I meant to put -6 in the f'(x), not f(x)

OpenStudy (anonymous):

f'(x) is the slope of the line tangent to the curve at x.

OpenStudy (anonymous):

so its -31?

OpenStudy (anonymous):

f(x)=x^2+5x f'(x)=2x+5 f'(-6)=2*(-6)+5=-12+5=-7

OpenStudy (anonymous):

oh i see. sorry i was confused. so -7 would be the slope then right

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

alright thanks!

OpenStudy (anonymous):

Welcome^^